Sweeten Company had no jobs in progress at the beginning of March and no beginning inventories. The company has two manufacturing departments--Molding and Fabrication. It started, completed, and sold only two jobs during March—Job P and Job Q. The following additional information is available for the company as a whole and for Jobs P and Q (all data and questions relate to the month of March):

Molding Fabrication Total
Estimated total machine-hours used 2,500 1,500 4,000
Estimated total fixed manufacturing overhead $ 14,250 $ 17,550 $ 31,800
Estimated variable manufacturing overhead per machine-hour $ 3.10 $ 3.90
Job P Job Q
Direct materials $ 30,000 $ 16,500
Direct labor cost $ 34,600 $ 14,300
Actual machine-hours used:
Molding 3,400 2,500
Fabrication 2,300 2,600
Total 5,700 5,100

Sweeten Company had no underapplied or overapplied manufacturing overhead costs during the month.

Required:

For questions 1-8, assume that Sweeten Company uses a plantwide predetermined overhead rate with machine-hours as the allocation base. For questions 9-15, assume that the company uses departmental predetermined overhead rates with machine-hours as the allocation base in both departments.

15. What was Sweeten Company’s cost of goods sold for March? (Do not round intermediate calculations.)
Answer #1

Calculate overhead rate :

Molding department = (14250/2500)+3.10 = $8.80 per machine hour

Fabrication department = (17550/1500)+3.90 = $15.60 per machine hour

13) calculate unit product Cost

Job P Job Q
Direct material 30000 16500
Direct labor 34600 14300
Overhead applied
Molding 3400*8.8 = 29920 2500*8.8 = 22000
Fabrication 2300*15.60 = 35880 2600*15.60 = 40560
Total product Cost 130400 93360

Cost of goods sold = 130400+93360 = $223760